# Real-Time Image Segmentation via Hybrid Convolutional-Transformer Architecture Search

### v2

@@ Abstract
Image segmentation is one of the most fundamental problems in computer vision and has drawn a lot of attention due to its vast applications in image understanding and autonomous driving. However, designing effective and efficient segmentation neural architectures is a labor-intensive process that may require numerous trials by human experts. In this paper, we address the challenge of integrating multi-head self-attention into high-resolution representation CNNs efficiently by leveraging architecture search. Manually replacing convolution layers with multi-head self-attention is non-trivial due to the costly overhead in memory to maintain high resolution. By contrast, we develop a multi-target multi-branch supernet method, which not only fully utilizes the advantages of high-resolution features but also finds the proper location for placing the multi-head self-attention module. Our search algorithm is optimized towards multiple objectives (e.g., latency and mIoU) and is capable of finding architectures on the Pareto frontier with an arbitrary number of branches in a single search. We further present a series of models via the Hybrid Convolutional-Transformer Architecture Search (HyCTAS) method that searches for the best hybrid combination of light-weight convolution layers and memory-efficient self-attention layers between branches from different resolutions and fuses them to high resolution for both efficiency and effectiveness. Extensive experiments demonstrate that HyCTAS outperforms previous methods in both semantic segmentation and panoptic segmentation tasks. Code and models are available at https://github.com/MarvinYu1995/HyCTAS.
@@
segmentation, high-resolution, neural architecture search, attention
journal: Journal of  Templates
@@ 1 Introduction
Over the past decade, deep neural networks have revolutionized segmentation performance [2, 3]. In particular, high-resolution representation architectures (e.g., HRNet [4, 5]) have significantly boosted accuracy by preserving spatial details throughout the network. However, these improvements often come at the cost of high computational overhead. Manually designing a segmentation framework that retains fine spatial features, incorporates global context (e.g., via attention mechanisms), and still runs in real time is notoriously difficult.
Table 1: Comparisons with state-of-the-art search methods [6, 7, 8, 9, 10]. In these comparisons, $\rho$ and $\tau$ represent the Pearson Correlation Coefficient and the Kendall Rank Correlation Coefficient of the true accuracy (by training each of them from scratch) and estimated accuracy (predicted by the super-network). Memory Cost reports the GPU memory consumption during the search process, GPU Days presents the searching cost, and Nums reports the number of the discovered architectures.
Method GPU Days Memory Cost Nums $\bm{\rho\uparrow}$ $\tau\uparrow$
Auto-DeepLab [6] 3 Total supernet 1 0.31 0.21
FasterSeg [7] 2 Total supernet 1 0.35 0.25
SparseMask [8] 4.2 Total supernet 1 0.49 0.38
DPC [9] 2600 One subnet 1 0.46 0.37
Fast-NAS [10] 8 One subnet 1 0.42 0.33
HyCTAS(Ours) 5 One subnet multiple 0.79 0.58
@@
In this paper, we propose a Hybrid Convolutional-Transformer Architecture Search (HyCTAS) framework to automatically discover powerful yet efficient segmentation networks. Rather than performing a manual, trial-and-error insertion of self-attention layers into a high-resolution CNN, HyCTAS leverages a multi-branch supernet search space and a genetic-based multi-objective optimization strategy. By jointly optimizing for latency and accuracy (mIoU), HyCTAS identifies an optimal “mix” of convolution and attention at multiple scales and resolutions. As a result, we obtain a family of architectures that lie on the Pareto frontier of speed and performance.
@@
    Novel multi-branch search framework: We encode parallel CNN branches at multiple resolutions, each of which may contain either lightweight convolution or memory-efficient attention. This unified supernet preserves high-resolution feature streams while optimizing for real-time inference.
@@
    Flexible multi-objective optimization: We employ a genetic-based algorithm to search for efficient architectures that yield a range of trade-offs between speed (latency) and segmentation accuracy (mIoU). Importantly, we can discover multiple Pareto-optimal architectures in a single search.
@@ 2.1 Efficient Networks for Image Segmentation
Semantic segmentation has long benefited from fully convolutional networks [13], which effectively adapt classification backbones to dense prediction tasks. Subsequent refinements, such as atrous/dilated convolutions in DeepLab [14] and multi-branch high-resolution designs in HRNet [4, 5], steadily pushed performance. Meanwhile, real-time or low-complexity methods such as ICNet [15], BiSeNet [16], and Fast-SCNN [17] introduced architectural optimizations (e.g., factorized convolutions, multi-resolution branches) to reduce latency. In more recent works, lightweight backbone designs [18, 19] target embedded devices and further illustrate the trade-off between efficiency and accuracy.
@@
Table 2: The difference between a typical operation and the proposed efficient operation. FLOPs are measured using an input of size 1$\times$256$\times$32$\times$64. The receptive field (RF) is relatively compared with the standard convolution.
Operation Type FLOPs (G) Params (M) RF Scale-to-Input
Convolution $1.2$ $0.59$ $1$ Linear
Light-weight Convolution $0.4$ $0.66$ $2$ Linear
Transformer 4.3 12.62 $Max$ Quadratic
Memory-efficient Self-attention $0.6$ $0.81$ $Max$ Quadratic
@@ 2.2 Neural Architecture Search
Recent works have extended NAS to semantic segmentation [23, 6, 7, 24], but many still rely on a single-path or single-resolution paradigm. Our HyCTAS goes further by accommodating multiple branches (with varying spatial resolutions) within one unified supernet. Moreover, our multi-objective search can discover a set of Pareto-optimal architectures—each balancing accuracy and speed differently—in a single pass.
@@ 2.3 Transformer in Vision Tasks
Transformers [25] originally showed remarkable success in NLP, and have recently gained popularity in vision tasks. Fully attentional backbones like Vision Transformer [11] and SETR [12] capture global context effectively but incur high memory costs, especially for large input resolutions. Hybrid strategies like DETR [26] combine CNNs with transformers for detection and segmentation.
Figure 1: An illustration of our HyCTAS search space. It preserves the multi-resolution and multi-branch properties of HRNet by introducing two types of searchable components: cells and nodes, which extend HRNet with multi-module fusion ability.
We follow the hybrid philosophy, blending lightweight convolutions with a memory-efficient self-attention module. Crucially, we do not merely handcraft these combinations; our search algorithm determines the where and when to use attention across multiple scales, striking a better balance between accuracy and runtime efficiency.
@@ 3.1 Overall Workflow
    Architecture Selection: The search yields a set of Pareto-optimal architectures, each offering a distinct trade-off (e.g., ultra-fast inference vs. higher accuracy). We finally retrain the chosen models from scratch to confirm their performance.
@@ 3.3 Combining Multiple Modules
Memory-efficient Self-attention Module: Instead of directly borrowing typical transformer operations like previous works [12, 28], we only employ the self-attention layer [25] without the FFN to capture the global context while maximizing efficiency. We then use $1\times 1$ convolutions for reducing and then increasing dimensions, leaving the self-attention module as a bottleneck with smaller than input/output channels. This largely reduces the memory usage required in the self-attention layer. Meanwhile, we add an extra bypassed $1\times 1$ convolution module for preserving uncompressed residue. Our memory-efficient self-attention module operation can utilize the largest receptive field while keeping relatively small FLOPs and memory footprint.
@@
3.4 Searching with Multi-objective
@@
${Maximize\;\;\;\;y=f(\alpha)=({f_{1}}(\alpha),{f_{2}}(\alpha),...,{f_{k}}(\alpha))}\\ {S.t.\;\;\;\;\;\;\;\;\;\;\;e(x)=({e_{1}}(\alpha),{e_{2}}(\alpha),...,{e_{m}}(\alpha))\leq 0}\\ {{\rm{among\;\;\;\;\;\;\;\;\;\;\;x}}=({\alpha_{1}},{\alpha_{2}},...,{\alpha_{n}})\in\mathcal{A}}\\ {\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;y=({y_{1}},{y_{2}},...,{y_{k}})\in Y}\end{array}$ (7)
@@
The detailed search algorithm is presented in Algorithm 1. One-hot encoding is employed to encode the sub-network $\alpha$. When the sampling is over, we will decode the sub-network to evaluate its corresponding mIoU and latency. We solve this multi-objective optimization (MOP) problem under the NSGA-II [30] framework.
Algorithm 1 Multi-objective Search.
  Input: supernetwork $\mathcal{N}$ with weights $\mathcal{W}$, the number of branches $b$, the number of generations $N$, population size $n$, validation dataset $D_{val}$, objective weights $w$.
  Output: $K$ individual architectures on the Pareto front.
  Uniformly generate the populations $P_{0}$ and $Q_{0}$ until each has $n$ individuals architectures.
  for $i=0$ to $N-1$ do
     $R_{i}=P_{i}\cup Q_{i}$
     $F,S=\text{Non-Dominated-Sorting}(R_{i})$
     Pick $n$ individual architectures to form $P_{i+1}$ by ranks and the crowding distance weighted by $w$.
     $Q_{i+1}=\emptyset$
     while $size(Q_{i+1})<n$ do
        $M=\text{Tournament-Selection}(P_{i+1})$
        $q_{i+1}=\text{Crossover}(M)$
        if $Latency(q_{i+1})>Latency({S_{max}})$ then
           continue
        end if
        Evaluate model $q_{i+1}$ with $S$ on $D_{val}$
        if $mIoU_{(}q_{i+1})>mIoU_{min}$ then
           Add $q_{i+1}$ to $Q_{i+1}$
        end if
     end while
  end for
  Select the best performing architecture from Pareto-front from $P_{N}$ on $D_{val}$
@@ 4 Experiments
Figure 3: The searched architectures under different objectives on the Cityscapes validation set. Our proposed HyCTAS approach can search for multiple efficient architectures with arbitrary branches in a single run. The color from light to dark indicates that our search results are getting better and better as the search progresses. The solid line represents architectures on the Pareto frontier. Best viewed in color.
Figure 4: An illustration of our searched architectures with multiple branches.

### v3

# HyCTAS: Multi-Objective Hybrid Convolution-Transformer Architecture Search for Real-Time Image Segmentation

@@ Abstract
Real-time image segmentation demands architectures that preserve fine spatial detail while capturing global context under tight latency and memory budgets. Image segmentation is one of the most fundamental problems in computer vision and has drawn a lot of attention due to its vast applications in image understanding and autonomous driving. However, designing effective and efficient segmentation neural architectures is a labor-intensive process that may require numerous trials by human experts. In this paper, we address the challenge of integrating multi-head self-attention into high-resolution representation CNNs efficiently by leveraging architecture search. Manually replacing convolution layers with multi-head self-attention is non-trivial due to the costly overhead in memory to maintain high resolution. By contrast, we develop a multi-target multi-branch supernet method, which not only fully utilizes the advantages of high-resolution features but also finds the proper location for placing the multi-head self-attention module. Our search algorithm is optimized towards multiple objectives (e.g., latency and mIoU) and is capable of finding architectures on the approximate Pareto front with an arbitrary number of branches in a single search. We further present a series of models via the Hybrid Convolutional-Transformer Architecture Search (HyCTAS) method that searches for the best hybrid combination of lightweight convolution layers and memory-efficient self-attention layers between branches from different resolutions and fuses them at high resolution for both efficiency and effectiveness. On Cityscapes, ADE20K, and COCO, HyCTAS discovers competitive real-time models without ImageNet pretraining, delivering strong accuracy and latency trade-offs. Code and models are available at https://github.com/MarvinYu1995/HyCTAS.
@@
real-time segmentation, high-resolution networks, multi-objective neural architecture search, memory-efficient self-attention, lightweight convolution
journal: Neurocomputing
@@ 1 Introduction
Over the past decade, deep neural networks have revolutionized segmentation performance [2, 3]. In particular, high-resolution representation architectures (e.g. HRNet [4]) have significantly boosted accuracy by preserving spatial details throughout the network. However, these improvements often come at the cost of high computational overhead. Manually designing a segmentation framework that retains fine spatial features, incorporates global context (e.g. via attention mechanisms), and still runs in real time is notoriously difficult. Recent surveys underline this tension between accuracy and efficiency in modern segmentation systems [5, 6].
Despite the prevalence of hybrid CNN–Transformer designs, many hybrid NAS frameworks restrict the search to single-path or single-resolution spaces and yield a single model per run, which makes it difficult to retain high-resolution streams and to cover diverse latency budgets. HyCTAS addresses this by enabling a multi-branch, multi-resolution search that returns an approximate Pareto front of architectures in a single run, explicitly balancing mIoU and latency.
In this paper, multi-branch refers to parallel feature streams at different spatial resolutions that are maintained throughout the network (in the spirit of HRNet), and multi-objective refers to simultaneous optimization of multiple objectives (e.g., mIoU versus latency/FLOPs/params) under a multi-objective formulation. We use approximate Pareto front/non-dominated set to denote the set of architectures not dominated by any other evaluated candidates in our search space [7].
The remainder of the paper is organized as follows: Section 2 reviews recent segmentation and NAS works. Section 3 introduces the HyCTAS search space and algorithm. Section 4 reports experiments across Cityscapes, ADE20K, and COCO with ablations, followed by limitations and conclusions.
@@
In this paper, we propose a Hybrid Convolutional-Transformer Architecture Search (HyCTAS) framework to automatically discover powerful yet efficient segmentation networks. Rather than performing a manual, trial-and-error insertion of self-attention layers into a high-resolution CNN, HyCTAS leverages a multi-branch supernet search space and an evolutionary computation-based multi-objective optimization strategy (NSGA-II). By jointly optimizing for latency and accuracy (mIoU), HyCTAS identifies an optimal “mix” of convolution and attention at multiple scales and resolutions. As a result, we obtain a family of architectures that lie on the approximate Pareto front of speed and performance.
@@
    Multi-branch search framework: We encode parallel CNN branches at multiple resolutions, each of which may contain either lightweight convolution or memory-efficient attention. This unified supernet preserves high-resolution feature streams while optimizing for real-time inference.
@@
    Flexible multi-objective optimization: We adopt an evolutionary computation-based approach (NSGA-II) to search for efficient architectures that yield a range of trade-offs between speed (latency) and segmentation accuracy (mIoU). We explicitly return multiple candidates from the non-dominated set in a single search.
@@
    Practical contribution focus: The contribution lies primarily in the NAS strategy and search space design for hybrid high-resolution networks rather than in new fusion operators; the lightweight convolution and memory-efficient attention are modest adaptations chosen to make the search space effective and hardware-friendly.
  • 4.
@@ 2.1 Efficient Networks for Image Segmentation
Semantic segmentation has long benefited from fully convolutional networks [10], which effectively adapt classification backbones to dense prediction tasks. Subsequent refinements, such as atrous/dilated convolutions in DeepLab [11] and multi-branch high-resolution designs in HRNet [4], steadily pushed performance. Meanwhile, real-time or low-complexity methods such as ICNet [12], BiSeNet [13], and Fast-SCNN [14] introduced architectural optimizations (e.g., factorized convolutions, multi-resolution branches) to reduce latency. In more recent works, lightweight backbone designs [15, 16] target embedded devices and further illustrate the trade-off between efficiency and accuracy. Among recent lightweight designs, LRDNet adopts a compact encoder–decoder with a refined dual-attention module to enable real-time segmentation [17].
Recent hybrid CNN–Transformer segmenters continue to improve efficiency by combining local and global modeling. Examples include SegFormer [18] and RTLinearFormer [19], which emphasize lightweight attention for mobile/real-time settings. Our work differs by searching the hybrid high-resolution design rather than handcrafting a single architecture. Extensive surveys [20] and the emergence of recent lightweight and hybrid architectures [16, 18] underscore the ongoing advancements since 2020.
Recent advances in efficient segmentation include HSNet-L [21], which uses hierarchical similarity networks for few-shot segmentation; AFFormer-B [22], which employs adaptive frequency Fourier transforms for efficient attention; and Vim-S [23], which leverages bidirectional state space models for linear-time sequence modeling. These methods demonstrate diverse approaches to balancing accuracy and efficiency in modern segmentation tasks.
@@ 2.2 Neural Architecture Search
Recent works have extended NAS to semantic segmentation [26, 27, 28, 29], but many still rely on a single-path or single-resolution paradigm. Comprehensive surveys [20] and recent lightweight and hybrid designs [16, 18] highlight continuing progress beyond 2020. Our HyCTAS goes further by accommodating multiple branches (with varying spatial resolutions) within one unified supernet. Moreover, our multi-objective search can discover a set of non-dominated architectures—each balancing accuracy and speed differently—in a single pass.
Multi-objective NAS. Beyond single-objective accuracy, multi-objective formulations jointly consider latency/FLOPs/parameters. We adopt NSGA-II [7] for its strong performance and simplicity, and we follow the terminology recommended in recent discussions of multi-objective NAS benchmarking (e.g., reporting an approximation front/non-dominated set rather than claiming true Pareto optimality on vast spaces). See [7] and the broader benchmarking perspective summarized in contemporary surveys [20].
@@ 2.3 Transformer in Vision Tasks
Transformers [30] originally showed remarkable success in NLP, and have recently gained popularity in vision tasks. Fully attentional backbones like Vision Transformer [8] and SETR [9] capture global context effectively but incur high memory costs, especially for large input resolutions. Hybrid strategies like DETR [31] combine CNNs with transformers for detection and segmentation. An efficient hybrid example is RTLinearFormer, which couples linear attentions with a dual-resolution backbone to reach real-time speed on urban-scene datasets [19].
Transformers excel at modeling long-range dependencies but incur quadratic cost in memory and computation at high resolutions [30, 8, 9]. We therefore blend lightweight convolutions with a memory-efficient self-attention module and let the search decide where/when to apply attention across scales, striking a better accuracy–runtime balance than manual designs.
@@ 3.1 Overall Workflow
    Architecture Selection: The search yields a set of non-dominated (approximate Pareto-optimal) architectures, each offering a distinct trade-off (e.g., ultra-fast inference vs. higher accuracy). We finally retrain the chosen models from scratch to confirm their performance.
For clarity, we use multi-branch to denote parallel streams operating at different resolutions that are maintained throughout the network, and multi-objective to denote our multi-objective optimization over both accuracy (mIoU) and latency. The search returns an approximate Pareto front (non-dominated set) of architectures.
Figure 1: An illustration of our HyCTAS search space. It preserves the multi-resolution and multi-branch properties of HRNet by introducing two types of searchable components: cells and nodes, which extend HRNet with multi-module fusion ability. Blue circles denote cells (operator selection; lightweight conv or memory-efficient self-attention), yellow circles denote nodes (fusion/skip gates controlling inter-branch connections). Solid arrows indicate active connections for a sampled sub-network; dashed arrows indicate optional connections available in the supernet that may or may not be selected during search. Upsampled features are concatenated and fused by a $3\times 3$ conv.
@@ 3.2 Revisiting HRNet Architecture
We use three scales (1/8, 1/16, 1/32) rather than four as in HRNet because keeping a 1/4 branch with attention significantly increases memory and latency for high-resolution inputs. Instead, we preserve 1/4 information in the stem and early fusion, and we allow the search to place attention at shallower stages when operating with fewer branches. This choice improves search tractability and real-time inference while retaining accuracy.
[…]
@@
Regarding the search space, cells choose between lightweight convolution and memory-efficient self-attention (Section 3.3), while nodes decide whether and how to fuse incoming branches at each resolution. The branch count is variable (1–3), and depths/widths per branch are adjustable. This compositional design yields a large combinatorial space yet remains structured enough for efficient search.
@@ 3.3 Combining Multiple Modules
[…]
Operation Type FLOPs (G) Params (M) RF Scale-to-Input
Convolution $1.2$ $0.59$ 11 Linear
Lightweight Convolution $0.4$ $0.66$ 22 Linear
Transformer 4.3 12.62 MaxMax Quadratic
Memory-efficient Self-attention $0.6$ $0.81$ MaxMax Quadratic
Why can the lightweight convolution have slightly more parameters than vanilla conv in Table 1? The parallel $1\times 1$ branch used to preserve high-resolution details adds a small number of parameters; in return, it reduces information loss from resizing and improves accuracy at similar or lower FLOPs. Our search can also choose it less frequently when strict parameter budgets are prioritized.
Memory-efficient Self-attention Module: Instead of directly borrowing typical transformer operations like previous works [9, 33], we only employ the self-attention layer [30] without the feed-forward network (FFN) to capture the global context while maximizing efficiency. We then use $1\times 1$ convolutions for reducing and then increasing dimensions, leaving the self-attention module as a bottleneck with smaller than input/output channels. This largely reduces the memory usage required in the self-attention layer. Meanwhile, we add an extra bypassed $1\times 1$ convolution module for preserving uncompressed residue. Our memory-efficient self-attention module operation can utilize the largest receptive field while keeping relatively small FLOPs and memory footprint.
In relation to linear/efficient attention, our module keeps exact (quadratic) dot-product attention but applies an aggressive channel bottleneck around attention to lower the memory/time cost. This differs from kernelized or low-rank ”linear” attention variants that approximate attention to achieve $\mathcal{O}(N)$ complexity; we chose exact attention with a bottleneck to maintain accuracy while controlling cost in high-resolution settings.
@@
We clarify that we do not claim the above modules as architectural novelties per se; rather, they are pragmatic choices that keep the search space expressive yet tractable in memory and latency, aligning with the main contribution of HyCTAS (the multi-objective NAS over multi-branch high-resolution backbones).
@@
3.4 Multi-objective Search
@@
[…]
@@
We encode candidates one-hot at the cell/node level and use tournament selection, one-point crossover at the branch-level topology, and bit-flip mutation over cell/operator choices with per-variable probability $1/n_{\text{var}}$, while respecting the constraints above. Non-dominated sorting and crowding distance then guide selection toward an approximate Pareto front (non-dominated set).
The detailed search algorithm is presented in Algorithm 1. One-hot encoding is employed to encode the sub-network $\alpha$. When the sampling is over, we will decode the sub-network to evaluate its corresponding mIoU and latency. We solve this multi-objective optimization (MOP) problem under the NSGA-II [7] framework. NSGA-II configuration. We use a seeded population of size $40$ for $20$ generations. Crossover is one-point at the branch-level topology with probability $p_{c}{=}0.9$; mutation is bit-flip over categorical genes with per-variable rate $1/n_{\text{var}}$. Duplicate architectures are eliminated, and selection uses non-dominated sorting with crowding-distance tournament. Searches are run independently for 1-, 2-, and 3-branch settings, yielding $40\times 20{=}800$ evaluations per setting and $2400$ in total.
Algorithm 1 Multi-objective Search
1:supernetwork $\mathcal{N}$ with weights $\mathcal{W}$, the number of branches $b$, the number of generations NN, population size $n$, validation dataset $D_{val}$, latency cap $S_{\max}$, minimum accuracy $\mathrm{mIoU}_{\min}$
2:$K$ architectures from the non-dominated set (approximate Pareto front)
3:Uniformly generate the populations $P_{0}$ and $Q_{0}$ until each has $n$ individual architectures.
4:for $i=0$ to N1N-1 do
5:  $R_{i}=P_{i}\cup Q_{i}$
6:  $F=\text{Non-Dominated-Sorting}(R_{i})$  # fronts and ranks
7:  Pick $n$ individual architectures to form $P_{i+1}$ according to the ranks and crowding distance.
8:  $Q_{i+1}=\varnothing$
9:  while $\text{size}(Q_{i+1})<n$ do
10:   $M=\text{Tournament-Selection}(P_{i+1})$
11:   $q_{i+1}=\text{Crossover}(M)$
12:   $q_{i+1}\leftarrow\text{Mutate}(q_{i+1})$
13:   if $\mathrm{Latency}(q_{i+1})>S_{\max}$ then
14:     continue
15:   end if
16:   $\mathrm{mIoU}(q_{i+1})\leftarrow\mathrm{Forward}(\mathcal{N}_{a};\mathcal{W})$ $\triangleright$ uses $\mathcal{W}$
17:   $\mathrm{Latency}(q_{i+1})\leftarrow\mathrm{TimeForward}(\mathcal{N}_{a};\mathcal{W})$ $\triangleright$ uses $\mathcal{W}$
18:   if $\mathrm{mIoU}_{(q_{i+1})}>\mathrm{mIoU}_{\mathrm{min}}$ then
19:     Add $q_{i+1}$ to $Q_{i+1}$
20:   end if
21:  end while
22:end for
23:Let $\mathcal{P}$ be the first front in $F$; return the top-$K$ by crowding distance from $\mathcal{P}$ for retraining.
For retraining, we take the returned $K$ architectures and train them from scratch with the schedule in Section 4; we then report final results on the validation/test sets using these retrained models.
@@ 4 Experiments
[…]
Figure 3: The searched architectures under different objectives on the Cityscapes validation set. Our proposed HyCTAS approach can search for multiple efficient architectures with arbitrary branches in a single run. The color from light to dark indicates that our search results are getting better and better as the search progresses. The solid line represents architectures on the approximate Pareto front (non-dominated set). Best viewed in color.